# Quest: Quality-Aware Metropolis-Hastings Sampling for Machine Translation

### v1

@@ 3.3 Connections to Reinforcement Learning with Human Feedback
$\alpha_{\beta}(y,y^{t})=\max\left\{1,\exp\left(\frac{r\left(y,x\right)-r\left(y^{t},x\right)}{\beta}\right)\frac{q(i|n)}{q(i|n^{t})}\right\}.$ (13)
The full derivation is in Appendix A. Using this approach, we can align language model generations without access to the model weights, log probabilities, or RL. We provide a preliminary experimental comparison using the two target distributions (Eq. 3 and Eq. 12) using Quest in Appendix D.
@@ Appendix A Connection to RLHF derivation
$\alpha_{\beta}(y,y^{t})=\max\left\{1,\exp\left(\frac{r\left(x,y\right)-r\left(x,y^{t}\right)}{\beta}\right)\frac{q(i|n)}{q(i|n^{t})}\right\}.$ (17)
@@ C.1 Toy problem: Validating the Approach
C.2 QE Results
We report the quality as measured by the metric optimized for, i.e., CometKiwi-XL in Figure 4. Quest results in higher quality across the board compared to ancestral sampling.
@@ C.2 QE Results
C.3 MCMC results in better outputs for shorter segments.
@@
C.4 Accepted proposals are skewed towards the end of the sentence.
@@
Appendix D Comparing RLHF-QUEST vs QUEST

### v2

@@ 1 Introduction
Figure 1: Quest samples an index from the current translation ($y^{t}$), removes all elements to the right of the index, generates a new continuation, and then uses the Metropolis-Hastings acceptance criterion to decide whether to accept or reject the resulting new translation. The process continues for a fixed number of $T$ iterations.
@@ 3.3 Connections to Reinforcement Learning with Human Feedback
$\alpha_{\beta}(y,y^{t})=\min\left\{1,\exp\left(\frac{r\left(y,x\right)-r\left(y^{t},x\right)}{\beta}\right)\frac{q(i|n)}{q(i|n^{t})}\right\}.$ (13)
The full derivation is in Appendix A. Using this approach, we can align language model generations without access to the model weights, log probabilities, or RL. We provide a preliminary experimental comparison using the two target distributions (Eq. 3 and Eq. 12) using Quest in Appendix C.4.
@@ Appendix A Connection to RLHF derivation
$\alpha_{\beta}(y,y^{t})=\min\left\{1,\exp\left(\frac{r\left(x,y\right)-r\left(x,y^{t}\right)}{\beta}\right)\frac{q(i|n)}{q(i|n^{t})}\right\}.$ (17)
@@ C.1 Toy problem: Validating the Approach
C.2 MCMC results in better outputs for shorter segments.
@@
C.3 Accepted proposals are skewed towards the end of the sentence.
@@
C.4 Comparing RLHF-QUEST vs QUEST
@@
C.5 QE Results
We report the quality as measured by the metric used in sampling, i.e., CometKiwi-XL in Figure 7. Quest results in higher quality across the board compared to ancestral sampling.
Figure 8: Average quality (xComet-XL) vs. diversity (Pairwise-BLEU) on additional LPs from WMT23 and WMT22. Different points represent different hyperparameter values.
C.6 Additional Language Pairs
[…]
C.7 COMET22 Results
We report the quality as measured by Comet22 (Rei et al., 2022) in Figure 9. Quest results in higher quality across the board compared to ancestral sampling.
[…]